OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 6x + 5y = 61 -5x + 5y = -5

OpenStudy (anonymous):

eq 1 minus eq 2 11x=66 x=6 6(6)+5y=61 5y=61-36 5y=25 y=5 Answer: (6,5)